OpenStudy (anonymous):

@ganeshie8 Jason has a red, a white, a blue, and a black shirt in his closet. He has dark blue, light blue, and black jeans. How many different combinations of shirt and jeans can Jason make?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

i think it 12

ganeshie8 (ganeshie8):

how many `shirts` Jason has ?

ganeshie8 (ganeshie8):

how many `jeans` Jason has ?

OpenStudy (anonymous):

ok sorry he has 4 shirts

OpenStudy (anonymous):

and three jeans

ganeshie8 (ganeshie8):

12 is right !

OpenStudy (anonymous):

wait A bag contains 6 green marbles, 18 red marbles, 2 white marbles, and 4 black marbles. What is the probability of drawing a black marble out of the bag? A.2/15 B.1/30 C.1/3 D.1/4

ganeshie8 (ganeshie8):

number of `black` marbles = ? number of `total` marbles = ?

OpenStudy (anonymous):

black marbles=4 total marbles=30

ganeshie8 (ganeshie8):

yes ! probability = (number of favorable outcomes) / (total number of outcomes)

ganeshie8 (ganeshie8):

just take the ratio

OpenStudy (anonymous):

B

ganeshie8 (ganeshie8):

No

OpenStudy (anonymous):

A?

ganeshie8 (ganeshie8):

probability of drawing a black marble = 4/30 = 2/15